OpenStudy (anonymous):

How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 31/2 yards long, and 5 pieces, each 21/4 yards long, are removed? I need help with the fraction part, answers have to be in fraction form, not decimal. Thanks!

OpenStudy (anonymous):

Well, you can just convert to decimals to make it easier to work with then convert back to fractions at the end.

OpenStudy (anonymous):

But if you want to work with fractions, just change 31/2 to quarters by multiplying top and bottom by 2.

OpenStudy (anonymous):

Then you add up all the pieces and subtract it from the total, 24 yards (you'll need to convert that into quarters to by multiplying top and bottom by 4)

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Um I know that 3 * 3 ½ is how I am supposed to do this, when I do I get 9 3/2 and for the second part I know that 5 * 2 ¼ is 10 5/4.

OpenStudy (anonymous):

Rule 1: Convert mixed numbers to improper fractions

OpenStudy (anonymous):

Wait. Is the question asking about 3 1/2 and 2 1/4, not 31/2 and 21/4?

OpenStudy (anonymous):

So 3 1/2=7/2 and 2 1/4 = 9/4

OpenStudy (anonymous):

In that case, yes, you would have to convert to improper fractions, like Tajar said- just multiply the number at the side by the denominator and add it to the numerator.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Rule #2: Translate the word problem into an equation

OpenStudy (anonymous):

You've got 3 7/2 pieces and 5 9/4 pieces . Now, you just need to multiply those numbers together to get the number of pieces.

OpenStudy (anonymous):

Here: 24 - [(7/2 x 3)] + [(9/4 x 5)] = solve!

OpenStudy (anonymous):

Once you've multiplied them, you'll have to add them to get the total amount removed.

OpenStudy (anonymous):

Since you want the total amount that remains, you'll have to then subtract the amount removed from the original amount.

OpenStudy (anonymous):

That's Tajar's equation.

OpenStudy (anonymous):

Hint: before solving, convert 24 to 96/4 and 21/2 to 42/4

OpenStudy (anonymous):

3 * 3 ½ = 10 ½ yards 5 * 2 ¼ = 11 ¼ yards Added 21 ¾ yards 24 yards - 21 ¾ yards = 2 ¼ yards left. Is this correct?

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

I believe. You should get into the habit of not using mixed numbers, they're a pain.

OpenStudy (anonymous):

Yes they are.

OpenStudy (anonymous):

Final Rule: Convert any improper fractions in your FINAL answer back to MIXED NUMBERS.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (anonymous):

Yw :)